# The State and Fate of Summarization Datasets

### v1

@@ Abstract
Automatic summarization has consistently attracted attention, due to its versatility and wide application in various downstream tasks. Despite its popularity, we find that annotation efforts have largely been disjointed, and have lacked common terminology. Consequently, it is challenging to discover existing resources or identify coherent research directions. To address this, we survey a large body of work spanning 133 datasets in over 100 languages, creating a novel ontology covering sample properties, collection methods and distribution. With this ontology we make key observations, including the lack in accessible high-quality datasets for low-resource languages, and the field’s over-reliance on the news domain and on automatically collected distant supervision. Finally, we make available a web interface that allows users to interact and explore our ontology and dataset collection, [^1] https://github.com/edahanoam/Awesome-Summarization-Datasets as well as a template for summarization data card, which can be used to streamline future research into a more coherent body of work.
The State and Fate of Summarization Datasets
@@ 1 Introduction
In this work, we conduct the most comprehensive survey of summarization datasets, covering a wide collection of 133 datasets across 104 languages. In Section 3, we develop a comprehensive ontology geared towards modern summarization datasets, positioning each dataset along seven different scales, including information about the samples (language, language modality, domain and summarization shape) as well as collection methods (annotation efforts and sources of supervision), and the manner in which the dataset is distributed.
Figure 1: There is a continuous increase in the number of summarization datasets being published. However, we find that the field is missing standardization and common terminology.
In Section 4 we show that our proposed ontology surfaces five major trends and gaps in the field. First, we find that the lack of common terminology thus far often leads to inconsistency and to the omission of important details when presenting datasets, making comparison challenging. An important example of this inconsistency is the traditional abstractive-extractive distinction, which we argue have gradually shifted away from a binary trait to a more nuanced spectrum.
Second, we observe that the field is over-reliant on the news domain, which acts as a double-edged sword – being both easy to obtain and prone to low quality. This is especially evident for low-resource languages, where the vast majority of datasets are based on news articles.
@@
Fourth, we find that most of the annotations were not specifically created to serve as summaries.
Fifth, we observe that copyright concerns, especially in the news domain have been changing data distribution, with many datasets move from publishing the data towards publishing scripts which reconstruct it from URLs. We argue that this practice is problematic, as URLs may change over time. Furthermore, a few seminal summarization datasets that were published publicly have been either taken down or are considered legally questionable due to copyright violations.
In Section 5 we publish two resources to standardize and help organize the vast research in the field. First, we make our dataset collection publicly available through a web UI which allows users to search datasets matching different criteria according to our terminology, for example, retrieving all manually created datasets in a certain language. This UI serves a central location where researchers can discover datasets, as well as identify new research directions. We welcome community contributions to keep this resource updated with newly published research.
Second, inspired by work on model cards Mitchell et al. (2019), we propose a standard summarization data card which covers the different items in our ontology as well as important properties. This can benefit future dataset developers in making their work more accessible, easy to discover and comparable against other datasets.
We hope that our work helps chart the way for future work in the field, as well enable researchers from various disciplines in exploring the wide range of existing summarization datasets.
@@
  • We formulate an ontology for summarization datasets, covering various axes regrading the samples, their collection and distribution.
  • We survey a wide range of 133 datasets covering datasets in 104 languages, showing that our ontology is rich enough to surface five important trends and challenges for future work, including the lack of terminology, the subpar quality of datasets in low-resource languages, and more.
@@ 2 Dataset Collection
Venues.
Most datasets are drawn from papers published in top-tier NLP, ML and AI venues and workshops. [^2] Venues: ACL, NAACL, EMNLP, LREC, EACL, SIGIR, AAAI, COLING, IEEE, NeurIPS, CoNLL, and TACL. In order to expand the supported languages, we also include datasets published through shared tasks and in language-specific journals and workshops.
Filtering.
We manually verify that each of the datasets is available for use, and exclude papers describing currently inaccessible resources. We include papers that explicitly focused on creating summarization datasets or that focused on summarization models and produced data as a by-product.
Collected Data.
For each dataset, we gather information on its properties, collection methods and distribution. This methodology allowed categorizing the current resources options for researchers, analyzing emerging trends and identifying the gaps in the field.
In order to analyze the current state of summarization datasets, we gathered additional information not covered in the ontology: the scale of data by number of samples, the levels of abstraction of the summaries and the data quality assessment methods applied to the collected data.
3 Ontology of Summarization Datasets
We propose a categorization of summarization datasets across seven orthogonal axes, and provide an ontology for each, depicted in Figure 2, and elaborated below. This ontology helps contextualize the extensive research in summarization and identifies missing types of datasets which could be addressed in follow up work.
3.1 Language and Language Modality
Monolingual.
Multilingual.
These datasets can be thought of as a few monolingual datasets together, as they contain document-summary pairs in various languages.
Cross-lingual.
3.2 Domain
Domain is the area of knowledge from which the documents are sourced. We identified nine domains that encompass the vast majority of datasets: News, which relies on articles from online platforms; Dialogue, which includes sources such as meeting transcripts, movie scenes, interviews, and customer service interactions; Scientific papers, most commonly paired with their abstract as summary; Legal documents such as government bills and patent documents; Encyclopedia entries, primarily based on Wikipedia; Opinions and arguments including reviews, editorials, and debates; Social media texts, which are typically more freeform; Literature such as books and poetry; and Instructional content, like "how-to" guides.
@@ 3.3 Summary Shape
We categorize the shape of the summaries in the datasets into four broad types, shown in Table 1.
@@
These are summaries ranging from a few words (like titles) to a lengthy sentence, composed of several dozens of words. This shape is sometimes referred to as “extreme summarization” (Narayan et al., 2018) or “headline generation” (Napoles et al., 2012).
@@
These summaries are constructed from a few, possibly independent sentences. For example, they could be bullet points (e.g., from news articles).
@@
These datasets contain output texts composed of one or more paragraphs of coherent free text. Examples of this shape include abstracts of scientific papers and sub-titles of news articles.
@@ 3.4 Sources for Supervision and Annotation Efforts
We categorize datasets by how much manual annotation effort they required. On opposite ends are datasets that were gathered in an Automatic way, mainly through scraping web resources, and human annotated datasets — in which annotators wrote the gold summaries. We also find semi-automatic datasets — in which human annotators are involved in the process but only later in the pipeline. Datasets created using this method typically provide annotators with summary options that they can correct, filter, or accept.
We argue that an additional often overlooked dimension is the objective for which the summaries were originally created, marked as Supervision in Figure 2.
@@
These datasets utilize summaries that were organically created. For example, highlights in news articles  Cheng and Lapata (2016); Hermann et al. (2015) and educational summaries published along text to help students Ladhak et al. (2020b).
@@
Similarly to the Natural category, these datasets use pre-existing texts as annotations. However, these texts were not originally created with the intention of serving as summaries, such as abstracts of scientific papers Cohan et al. (2018) or titles and subheadings of news articles Napoles et al. (2012).
@@
This category features datasets in which the summaries were written due to the creators’ efforts and did not exist beforehand. It includes mostly human annotated datasets, but also synthetically created ones.
@@ 3.5 Availability
The distribution of datasets depends on how accessible they are to the research community. We identify three main forms of data availability.
@@
These datasets are published through a platform like Hugging Face.
@@
In this category, the data is not directly published; instead, a method to reconstruct it is provided. For example, a list of URLs to scrape as done in Grusky et al. (2018), or data preparation scripts as done in Kryściński et al. (2021).
@@
This availability type, which is most restricting, usually requires filling a request form or contacting the authors in order to receive the data.
@@ 4 The State of Summarization Datasets
Figure 3: Every dot on the graph corresponds to a dataset in our collection. Each dot’s position indicates the percentage of unique uni-grams in the summaries that are not found in the source texts. Only a minority of the datasets report evaluation of the properties of the summaries.
We thoroughly examine 133 summarization datasets under the lens of our ontology. Figure 2 presents the quantitative dataset distribution as well as example categorization of several datasets. Following, we discuss five major observations regarding the current state and challenges in summarization datasets.
4.1 There is a Lack in Common Terminology
While summarization is a versatile and expressive task applicable in many scenarios, we find that summarization datasets are presented using inconsistent terminology and often lacking details. This makes it difficult for researchers to identify available resources and to effectively compare and contrast them.
For example, determining the shape of the summaries (such as paragraph or highlights) often required to access the dataset and examine samples; understanding how the summaries were created frequently involved visiting the original website from which the data was scraped; while datasets often do not mention the language in which they were annotated, in contrast with recent emphasis on explicit language disclosure Bender (2019).
All of these highlight the necessity of a shared ontology expressive enough to capture currently important aspects in summarization datasets, and guide dataset developers in reporting their work.
@@ 4.1 There is a Lack in Common Terminology
An important case where terminology is lacking is the distinction between extractive summarization, where models are required to select sentences from the source document and abstractive summarization, where generative models can also produce novel text.
With the advent of powerful generative LLMs, our survey indicates that these distinctions are no longer useful, with many of the surveyed datasets not specifying whether their annotations were either extractive or abstractive.
Instead, we argue that abstractiveness is a spectrum rather than a binary trait. To support this, in Figure 3 we plot the novel n-grams ratio for datasets in our collection. This metric, which we find to be the most commonly used for measuring levels of abstractions, measures the percentage of n-grams appearing in the summaries but not in the document. Formally Narayan et al. (2018):
@@
On this scale, 0% can be achieved by restricting the summary to contain only vocabulary from the source document and thus lean towards extractive. While at the other end of the scale, 40% of the 1-grams in the summaries were novel and thus leaning towards abstractive. Other less common metrics include Coverage and Density Grusky et al. (2018) as well as Abstractivity (ABS; Bommasani and Cardie, 2020). All of which are based on the overlap between the summary and the source document.
@@
4.2 News is the Most Popular Domain, Despite its Major Shortcomings
Roughly 50% of the datasets we surveyed use news articles as source documents. This popularity may be due to the large number of online news platforms in various languages that allow for easy scraping. Moreover, many of these articles already include human-written summaries or distant supervision, in the form of highlights, titles, and subtitles. We argue that this wide availability and relative ease of obtaining distant supervision is in fact a double-edged sword, harming the quality of the gold summaries, while presenting a rather simple challenge for summarization models.
First, the news domain may present a simple instance of the summarization task. News articles typically follow “inverted pyramid writing style” — i.e., the most important information answering the 5 w’s (who? what? etc.) appears in the first few sentences (Pottker, 2003). Consequently, a summarization model which trivially outputs the first sentences of the news article can often achieve good results (Cohen et al., 2021).
Second, news summarization datasets are prone to be of low quality, because they often rely on automatic heuristics. Tejaswin et al. (2021) manually evaluated two datasets with one sentence summaries, one which uses the title as summary Rush et al. (2015), and another which uses a bold sentence in the beginning of the article Narayan et al. (2018). They found that about half of the samples are either incomplete or contain concepts or entities that do not appear in the rest of the document, and hence often cannot be deduced from it.
Beyond the news domain, the scientific domain, characterized by longer texts and complex terminology is popular Cohan et al. (2018), yet it remains predominantly English-centric. Official documents, such as those in the legal domain, patents, and government bills (e.g., Sharma et al., 2019; Kornilova and Eidelman, 2019; Huang et al., 2021), share similar complexity in that salient information is usually distributed throughout the document, and are more multilingual. We also find that dialogues are increasingly attracting the attention of the research community, with this domain being the second most popular, and that Wikipedia articles are a valuable resource due to their guidelines and API, as well as for cross-lingual research because they cover the same topics in multiple languages.
@@ 4.3 Datasets for Low-Resource Languages are Often Low Quality and Hard to Discover
The attention to performing NLP research on a diverse set of languages has been increasing in recent years (Joshi et al., 2020; Ruder, 2022). Evidently, while English is by far the most popular language in our survey, we cover resources for 104 different languages (see Appendix A for the complete list). We find that this encouraging versatility was achieved mainly through multi-lingual datasets, as 81 of the languages surveyed are supported solely through them. In Figure 4, we illustrate the growth in the number of languages with available resources over the years, emphasizing the dramatic rise following the introduction of multilingual resources.
However, while providing important diversity, we identify limitations in how multilingual summarization datasets are created and reported.
Discovering existing datasets in low-resource languages proves challenging.
Given that the majority of languages in our survey are available only within large multilingual corpora,
researchers must comb through all multilingual datasets in search for their language. This is challenging as there is currently no unified platform that houses all available datasets and allows to search by the desired language.
@@
To make matters worse, determining which languages are present in each multilingual dataset proved challenging, as we mentioned before, datasets often do not explicitly list the supported languages in the paper.
@@
4.4 Most Annotations were not Intended to be Used as Summaries
We found only 35 dedicated datasets, i.e., where the annotations were created specifically for the task and did not exist before. This is three times fewer than datasets that include existing annotations.
Not only that the vast majority of annotations available were gathered automatically, but datasets collected this way are an order of magnitude bigger: human annotated datasets were on average more than 10 times smaller in number of samples than automatically collected datasets and thus allow for training big models.
We categorized 43% of the datasets as Distant supervision, meaning the human authors did not explicitly intend to write a summary of the source document. This includes proxies of summaries like titles and subtitles in news articles, abstracts of scientific papers or descriptions of videos or interviews. Other approaches we consider as distant are social media metadata as summaries, as introduced in Newsroom Grusky et al. (2018) and beginning with summaries and searching for longer texts to support them Ghalandari et al. (2020); Saxena and Keller (2024).
We also find a mismatch between the datasets we create and the ones we use. In Figure 6, we show a clear trend over the years of collecting the paragraph shape. This might be considered a natural shape for summarization as it is also the most common for collecting data through human annotation in our collection. This comes in contrast to Highlights being very commonly used in research. This is evident from CNN-Daily Mail, perhaps the most popular summarization dataset at the time of writing, which uses this shape.

### v2

@@ Abstract
Automatic summarization has consistently attracted attention due to its versatility and wide application in various downstream tasks. Despite its popularity, we find that annotation efforts have largely been disjointed, and have lacked common terminology. Consequently, it is challenging to discover existing resources or identify coherent research directions. To address this, we survey a large body of work spanning 133 datasets in over 100 languages, creating a novel ontology covering sample properties, collection methods and distribution. With this ontology we make key observations, including the lack of accessible high-quality datasets for low-resource languages, and the field’s over-reliance on the news domain and on automatically collected distant supervision. Finally, we make available a web interface that allows users to interact and explore our ontology and dataset collection, [^1] https://github.com/edahanoam/Awesome-Summarization-Datasets, [^2] https://searchdatasets-fru5zwwfm2shmyrtna9gjh.streamlit.app/ as well as a template for a summarization data card, which can be used to streamline future research into a more coherent body of work.
The State and Fate of Summarization Datasets: A Survey
@@ 1 Introduction
In this work, we conduct the most comprehensive survey of summarization datasets to date, covering a wide collection of 133 datasets across 104 languages. In Section 3, we develop a comprehensive ontology geared towards modern summarization datasets, positioning each dataset along seven different scales, including information about the samples (language, language modality, domain and summarization shape) as well as collection methods (annotation efforts and sources of supervision), and the manner in which the dataset is distributed.
Figure 1: The number of summarization datasets being published is continuously increasing, yet the field lacks standardization and common terminology.
In Section 4 we show that our proposed ontology surfaces five major trends and gaps in the field. First, we find that the lack of common terminology thus far often leads to inconsistency and to the omission of important details when presenting datasets, making comparison challenging. An important example of this inconsistency is the traditional abstractive-extractive distinction, which we argue has gradually shifted away from a binary trait to a more nuanced spectrum.
Second, we observe that the field is over-reliant on the news domain, which acts as a double-edged sword – being easy to obtain while prone to low quality. This is especially evident for low-resource languages, where the vast majority of datasets are based on news articles.
@@
Fourth, we find that while the majority of annotations are gathered automatically, they are typically not originally created to serve as summaries, relying instead on distant supervision.
Fifth, we observe that copyright concerns, especially in the news domain, have been changing data distribution, with many datasets moving from publishing the data to publishing scripts that reconstruct it from URLs. Furthermore, a few seminal summarization datasets that were published publicly have been either taken down or are considered legally questionable due to copyright violations.
In Section 5 we publish two resources to standardize and help organize the vast research in the field. First, we make our dataset collection publicly available through a web UI which allows users to search datasets matching different criteria according to our terminology, for example, retrieving all manually created datasets in a certain language. This UI can help researchers discover datasets, as well as identify new research directions. We welcome community contributions to keep this resource updated with new research.
Second, inspired by work on model cards Mitchell et al. (2019), we propose a standard summarization data card which covers the different items in our ontology as well as additional important properties. This can benefit future dataset developers in making their work more accessible, easy to discover, and comparable against other datasets.
We hope that our work helps chart the way for future work in the field, as well as enable researchers from various disciplines in exploring the wide range of existing summarization datasets.
@@
  • We formulate an ontology for summarization datasets, covering various axes regarding the samples, their collection and distribution.
  • We survey a wide range of 133 datasets covering 104 languages, showing that our ontology is rich enough to surface five important trends and challenges, including the lack of terminology, the subpar quality of datasets in low-resource languages, and more.
@@ 2 Dataset Collection
Venues.
Most datasets are drawn from papers published in top-tier NLP, ML and AI venues and workshops. [^3] Venues: ACL, NAACL, EMNLP, LREC, EACL, SIGIR, AAAI, COLING, IEEE, NeurIPS, CoNLL, and TACL. We also include datasets published through shared tasks and in language-specific journals and workshops.
Filtering.
We manually verify that each dataset is available for use, and exclude papers describing currently inaccessible resources. We include papers that explicitly focused on creating summarization datasets or that focused on summarization models and produced data as a by-product.
Collected Data.
For each dataset, we gather information on its properties, collection methods and distribution. This methodology allows categorizing the current resources available, analyzing emerging trends and identifying the gaps in the field.
3 Ontology of Summarization Datasets
We propose a categorization of summarization datasets across seven orthogonal axes, and provide an ontology for each, depicted in Figure 2, and elaborated below. This ontology helps contextualize the extensive research in summarization and identifies missing types of datasets which can be addressed in follow up work.
3.1 Language and Language Modality
Monolingual.
Multi-lingual.
These datasets can be thought of as a few monolingual datasets joined together, as they contain document-summary pairs in various languages.
Cross-lingual.
3.2 Domain
We identified nine domains that encompass the vast majority of datasets: News, which relies on articles from online platforms; Dialogue, which includes sources such as meeting transcripts, movie scenes, interviews, and customer service interactions; Scientific papers, most commonly paired with their abstract as a summary; Legal documents such as government bills and patent documents; Encyclopedia entries, primarily based on Wikipedia; Opinions and arguments including reviews, editorials, and debates; Social media texts, which are typically more freeform; Literature such as books and poetry; and Instructional content, like "how-to" guides.
@@ 3.3 Summary Shape
We categorize the shape of the summaries into four broad types, exemplified in Table 1.
@@
These are summaries ranging from a few words (like titles) to a full sentence, composed of several dozens of words. This shape is sometimes referred to as “extreme summarization” (Narayan et al., 2018) or “headline generation” (Napoles et al., 2012).
@@
These summaries are constructed from a few, possibly independent sentences. For example, they could be bullet points from news articles.
@@
These datasets contain output texts composed of one or more paragraphs of coherent free text. Examples of this shape include abstracts of scientific papers and sub-headlines of news articles.
@@ 3.4 Sources for Supervision and Annotation Efforts
We categorize datasets by how much manual annotation effort they require. On opposite ends are datasets that were gathered in an Automatic way, mainly through scraping web resources, and Human Annotated datasets — in which annotators wrote the gold summaries. We also find Semi-Automatic datasets — in which human annotators are involved in the process in varying degrees. Datasets created using this method typically provide annotators with summary options that they can correct, filter, or accept.
An additional often overlooked dimension is the objective for which the summaries were originally created, marked as Supervision in Figure 2, and consisting of the following possible values:
@@
These datasets utilize summaries that were organically created. For example, highlights in news articles  Cheng and Lapata (2016); Hermann et al. (2015) and educational summaries published to help students Ladhak et al. (2020b).
@@
Similarly to the Natural category, these datasets use pre-existing texts as annotations. However, these texts were not originally created to serve as summaries, such as abstracts of scientific papers Cohan et al. (2018) or titles and sub-headlines of news articles Napoles et al. (2012).
@@
This category features datasets in which the summaries were written especially for the dataset. It includes mostly human annotated datasets, but also synthetically created ones.
@@ 3.5 Availability
The distribution of datasets denotes how accessible they are to the research community. We identify three main forms of data availability.
@@
These datasets are often published through a platform like Hugging Face.
@@
In this category, the data is not directly published; instead, a method to reconstruct it is provided. For example, a list of URLs as done in Grusky et al. (2018), or data preparation scripts, as in Kryściński et al. (2021).
@@
This availability type, which is most restricting, usually requires filling out a request form or contacting the authors in order to receive the data.
@@ 4 The State of Summarization Datasets
Figure 3: Every dot on the graph corresponds to a dataset in our collection. Each dot’s position indicates the percentage of unique uni-grams in the summaries that are not found in the source texts. Only a minority of the datasets report an evaluation of the properties of the summaries.
We thoroughly examine 133 summarization datasets under the lens of our ontology. Figure 2 presents the quantitative dataset distribution as well as categorization examples for several datasets. Following, we discuss five major observations regarding the current state and challenges in summarization datasets.
4.1 Lack in Common Terminology
While summarization is a versatile and expressive task applicable in many scenarios, we find that summarization datasets are presented using inconsistent terminology and often lacking details. This makes it difficult for researchers to identify available resources and to effectively compare them.
For example, determining the shape of the summaries (such as paragraph or highlights) often requires accessing the dataset and examining samples; understanding the supervision, how the summaries were created, frequently involved visiting the original website from which the data was scraped; while datasets often do not mention the language in which they were annotated, in contrast with recent emphasis on explicit language disclosure Bender (2019). In our collection, 39 papers failed to specify their languages, often assuming that English is the default language.
All of these highlight the necessity of a shared ontology expressive enough to capture important aspects in summarization datasets, and guide dataset developers in reporting their work.
@@ 4.1 Lack in Common Terminology
Another example where terminology is lacking is the distinction between extractive summarization, where models are required to select text from the source document and abstractive summarization, where generative models can produce novel text.
With the advent of powerful generative LLMs, our survey indicates that these distinctions are no longer useful, with many of the surveyed datasets not specifying if their annotations were extractive or abstractive.
Instead, we argue that abstractiveness is a spectrum rather than a binary trait. To support this, in Figure 3 we plot the novel n-grams ratio for datasets in our collection. This metric, which we find to be the most commonly used for measuring levels of abstractions, measures the percentage of n-grams appearing in the summaries but not in the document Narayan et al. (2018). Formally:
@@
On this scale, 0% can be achieved by restricting the summary to contain only vocabulary from the source document and thus lean towards extractive. While at the other end of the scale, 40% of the 1-grams in the summaries were novel and thus leaning towards abstractive. Other metrics for measuring abstractiveness include Coverage and Density Grusky et al. (2018) as well as Abstractivity (ABS; Bommasani and Cardie, 2020). All of which are based on the overlap between the summary and the source document.
4.2 News is the Most Popular Domain, Despite its Major Shortcomings
Roughly 50% of the datasets in our survey use news articles as source documents. This popularity may be due to the large number of online news platforms in various languages that allow for easy scraping. Moreover, many of these articles already include human-written summaries or distant supervision, in the form of highlights, titles, and sub-headlines. This wide availability and relative ease of obtaining distant supervision is in fact a double-edged sword, harming the quality of the gold summaries, while presenting a rather simple challenge for summarization models.
First, the news domain may present a simple instance of the summarization task. News articles typically follow an “inverted pyramid writing style” — i.e., the most important information answering the 5 w’s (who? what? etc.) appears in the first few sentences (Pottker, 2003). Consequently, a summarization model that trivially outputs the first sentences of a news article can often achieve good results (Cohen et al., 2021).
Second, news summarization datasets are prone to be of low quality, because they often rely on automatic heuristics. Tejaswin et al. (2021) manually evaluated two datasets with one sentence summaries, one that uses the title as a summary Rush et al. (2015), and another that uses a bold sentence at the beginning of the article Narayan et al. (2018). They found that about half of the samples are either incomplete or contain concepts or entities that do not appear in the rest of the document, and hence often cannot be deduced from it.
Beyond the news domain, the scientific domain, characterized by longer texts and complex terminology is popular Cohan et al. (2018), yet it remains predominantly English-centric. Official documents, such as those in the legal domain, patents, and government bills (e.g., Sharma et al., 2019; Kornilova and Eidelman, 2019; Huang et al., 2021), share similar complexity in that salient information is usually distributed throughout the document, and are more multilingual. We also find that dialogues are increasingly attracting the attention of the research community, with this domain being the second most popular, and that Wikipedia articles are a valuable resource due to their guidelines and API, and they are useful for cross-lingual research as they cover the same topics in multiple languages.
@@ 4.3 Datasets for Low-Resource Languages are Often Low Quality and Hard to Discover
The attention to multilingual NLP has been increasing in recent years (Joshi et al., 2020; Ruder, 2022). Evidently, while English is by far the most popular language in our survey, we cover resources for 104 different languages (see Appendix A for the complete list). We find that this encouraging versatility was achieved mainly through multi-lingual datasets, as 81 of the languages surveyed are supported solely through them. In Figure 4, we illustrate the growth in the number of languages with available resources over the years, emphasizing the dramatic rise following the introduction of multilingual resources.
However, we identify limitations in how multilingual datasets are created and reported.
Discovering datasets in low-resource languages proves challenging.
Given that the majority of languages in our survey are available only within large multilingual corpora, researchers must sift through all multilingual datasets to find their language. This is challenging as there is currently no unified platform that houses all available datasets and allows searching by the desired language.
@@
4.4 Annotations Were Often Not Intended to Be Used as Summaries
We categorize 43% of the datasets as Distant supervision, meaning human authors did not explicitly intend to write a summary of the source document. This is the most common supervision category. It includes proxies of summaries like titles and sub-headlines of news articles, abstracts of scientific papers or descriptions of videos and interviews. Other approaches we consider as distant are social media metadata as summaries, as introduced in Newsroom Grusky et al. (2018) and starting with summaries before searching for longer texts to support them Ghalandari et al. (2020); Saxena and Keller (2024).
On the other end of the supervision spectrum are dedicated datasets, where the annotations were created specifically for the dataset and did not exist before. We found only 35 dedicated datasets, three times fewer than datasets that include existing annotations.
In addition to the vast majority of annotations being gathered automatically, datasets collected this way are also an order of magnitude bigger: human-annotated datasets are on average more than 10 times smaller in number of samples than automatically collected datasets and thus allow for training big models.
We also find a mismatch between the datasets created and those used. In Figure 6, we show a clear trend over the years of collecting the paragraph shape. This might be considered a natural shape for summarization as it is also the most common for collecting data through human annotation in our collection. This comes in contrast to Highlights being very commonly used in research. This is evident from CNN-Daily Mail, perhaps the most popular summarization dataset at the time of writing, which uses this shape.